OpenStudy (anwar3341):

I need the answer for this question please !se a table of numerical values of f(x, y) for (x, y) near the origin to make a conjecture about the value of the limit of f(x, y) as (x,y) ->(0,0) f(x,y) =x^2*y^3+x^3*Y2-3/ (1-xy)

OpenStudy (irishboy123):

linearise it !

OpenStudy (anwar3341):

how can I linearise it ?!

OpenStudy (irishboy123):

well, currently, you want x and y to be zero. so your answer is zero.

OpenStudy (anwar3341):

there is no Zero in the multiple choices

OpenStudy (anwar3341):

3 or -1 or -3 or 1 or the value can not be determined

OpenStudy (anwar3341):

?

OpenStudy (irishboy123):

perhaps i am reading this the wrong way. and maybe, i think, it's this: \(f(x,y) =x^2y^3+x^3y^2-\dfrac{3}{1-xy} \) so then \(f(0,0) = - \dfrac{3}{1}\) IOW you don't need a table does that make any sense to you?? doesn't to me :(